Ammonium carbamate dissociates as : In a closed vessel containing ammonium carbamate in equilibrium with ammonia and carbon dioxide, ammonia is added such that partial pressure of NH3 now equals to the original total pressure. The ratio of total pressure to the original pressure now is ?

NH2COONH4(s)----2NH3(g)+CO2(g)

Grade:11

1 Answers

Arun
5 years ago
 after adding NH3 the equilibrium pressure of NH3 becomes p which is the original pressure.since Kp will be same for the above reaction.Therefore4p^3/27=p^2*xwhere x is the equilibrium partial pressure of CO2.x=4p/27Thus the total pressure=4p/27+p=31p/27The ratio is =31p/27/p=31/27